OpenStudy (anonymous):

AP Calculus: The radius of a sphere is measured to be 3 inches. If the measurement is correct to within + or - .01 inch, use differentials to estimate the possible error in the volume of the sphere.

OpenStudy (anonymous):

@Fall12-13 do you get how to do this?

OpenStudy (anonymous):

No idea

OpenStudy (anonymous):

@agent0smith do you know how to do this problem? :D

OpenStudy (agent0smith):

I can give it a shot

OpenStudy (anonymous):

btw answer is \[\pm .36\pi \]

OpenStudy (agent0smith):

Alright. You can probably also check the answer just by putting r=3.01 and r=2.99 inches into the equation for volume of a sphere, it should give a similar figure. Sphere volume: \[V=\frac{ 4 }{ 3 }\pi r ^{3}\] I think we just want to find dV/dr and then use dr as 0.01 inches and find dV... \[\frac{ dV }{ dr }=4 \pi r ^{2}\]

OpenStudy (agent0smith):

Yeah that's all we need to do, it's actually kinda easy. Now we have dV/dr we can treat it as a fraction with dr = 0.01 inches, and find dV which will be the error in volume. \[dV=4\pi r ^{2} dr\]

OpenStudy (agent0smith):

r = 3 inches and dr = +/- 0.01 inches, so substitute those in. For +0.01 you'll get dR = +0.36pi cubic inches, and for -0.01 you'll get dR = -0.36pi cubic inches. So dV = +/- 0.36pi cubic inches

OpenStudy (anonymous):

ah ok ty <333

OpenStudy (anonymous):

@agent0smith can you help me with another one please? :)

OpenStudy (agent0smith):

yeah probably